### v1

@@ 3.2 The update of $[X_{i}]$
where $[P_{i}^{k}]=(\mathbf{R}(\mathcal{L}^{k}))_{[(i)]}$, $[Q_{1}^{k}]=\left([\mathcal{G}^{k}]\times_{2}[X^{k}_{2}]\times_{3}[X^{k}_{3}]\right)_{[(1)]}$, $[Q_{2}^{k}]=\big{(}[\mathcal{G}^{k}]\times_{1}[X_{1}^{k+1}]$$\times_{3}[X^{k}_{3}]\big{)}_{[(2)]}$, $[Q_{3}^{k}]=\left([\mathcal{G}^{k}]\times_{1}[X^{k+1}_{1}]\times_{2}[X^{k+1}_{2}]\right)_{[(3)]}$, and parameter $\tilde{\alpha}_{X}=\frac{1}{1+\alpha_{X}}$.
@@
Lemma 3.1.
[17, 2] Given $A\in\mathbb{R}^{m\times n}$, $m\geq n$, consider the following Stiefel manifold projection problem
@@ 3.3 The update of $[\mathcal{G}]$
(11) $[\mathcal{G}^{k+1}]=\operatorname{prox}_{\tilde{\alpha}_{\mathcal{G}}\psi}\left([\mathcal{G}^{k}]-\tilde{\alpha}_{\mathcal{G}}([\mathcal{G}^{k}]-[\mathcal{O}^{k+1}])\right),$
where $[\mathcal{O}^{k+1}]=\mathbf{R}(\mathcal{L}^{k})\times_{1}[X_{1}^{k+1}]^{\top}\times_{2}[X_{2}^{k+1}]^{\top}\times_{3}[X_{3}^{k+1}]^{\top}$, and $\tilde{\alpha}_{\mathcal{G}}=\frac{1}{1+\alpha_{\mathcal{G}}}$.
@@ 3.4 The update of $\mathcal{L}$
After computing $[X_{i}^{k+1}]$ and $[\mathcal{G}^{k+1}]$, we can obtain the approximated low-rank group tensor, denoted as $[\mathcal{Y}^{k+1}]$, as follows
(12) $[\mathcal{Y}^{k+1}]=[\mathcal{G}^{k+1}]\times_{1}[X_{1}^{k+1}]\times_{2}[X_{2}^{k+1}]\times_{3}[X_{3}^{k+1}].$
Then $\mathcal{L}^{k+1}$ can be computed in a unique closed form as follows
(13) $\mathcal{L}^{k+1}=$ $\tilde{\delta}\mathcal{W}_{\mathbf{R}}^{-1}\odot\mathbf{R}^{\top}([\mathcal{Y}^{k+1}])+(1-\tilde{\delta})(\mathcal{D}-\mathcal{S}^{k+1}),$
where parameter $\tilde{\delta}=\frac{1}{1+\delta}$ and $\mathcal{W}_{\mathbf{R}}^{-1}$ denotes the component-wise inverse of $\mathcal{W}_{\mathbf{R}}$, i.e., the $(i_{1},i_{2},i_{3})$-th entry of $\mathcal{W}_{\mathbf{R}}^{-1}$ is equal to $1/(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}$.
@@ 4 Convergence Analysis of The P-BCD Algorithm
${0}\in\delta\mathcal{W}_{\mathbf{R}}\odot(\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D})+\gamma\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\partial\varphi(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\bar{\mathcal{S}}),$
@@
${0}\in\nabla_{[\mathcal{G}]}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})+\partial\psi([\bar{\mathcal{G}}]),$
${0}=\delta\mathcal{W}_{\mathbf{R}}\odot(\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D})+\nabla_{\mathcal{L}}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),$
@@
(15a) ${0}\in\delta(\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D})+\gamma\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\right)^{-1}\odot\partial\varphi(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\bar{\mathcal{S}}),$
@@
(15e) ${0}\in[\bar{\mathcal{G}}]-[\bar{\mathcal{O}}]+\partial\psi([\bar{\mathcal{G}}]),$
(15f) ${0}=\delta(\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D})+\bar{\mathcal{L}}-\mathcal{W}_{\mathbf{R}}^{-1}\odot\mathbf{R}^{\top}([\bar{\mathcal{Y}}]),$
@@
$\geq$ $\frac{\alpha_{\mathcal{S}}}{2}\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot(\mathcal{S}^{k+1}-\mathcal{S}^{k})\|^{2}_{F}$
$\geq$ $\frac{c_{1}^{2}\alpha_{\mathcal{S}}}{2}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F},$
where $c_{1}=\min\,(\sqrt{\mathcal{W}_{\mathbf{R}}})_{i_{1}i_{2}i_{3}}$.
Next, it follows from the update of $[X_{i}]$ and Lemma 3.1 that
@@
Combining the inequalities above, we obtain (16) with $\alpha=\min\{c_{1}^{2}\alpha_{\mathcal{S}},\alpha_{\mathcal{G}},\alpha_{X}\}$.
(ii) Since $[X_{i}^{k}]^{\top}[X_{i}^{k}]=[\mathbf{I}_{n_{i}}]$ for each $i=1,2,3$, we have the sequence $\{[X_{i}^{k}]\}$ is bounded. By (i), we have $\Phi(\mathcal{Z}^{k})\leq\Phi(\mathcal{Z}^{0})$. Also, we observe that $\Phi(\mathcal{Z}^{k})\geq\gamma c_{1}^{p}\varphi(\mathcal{S}^{k})+\psi([\mathcal{G}^{k}])\geq 0$. Since
@@
$\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}\leq$ $\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}+c_{2}\|[{X}_{1}^{k+1}]-[{X}_{1}^{k}]\|_{F}$
$+c_{2}\|[{X}_{2}^{k+1}]-[{X}_{2}^{k}]\|_{F}+c_{2}\|[{X}_{3}^{k+1}]-[{X}_{3}^{k}]\|_{F},$
where $c_{2}=\max_{k}\|[\mathcal{G}^{k}]\|_{F}<\infty$ according to assertion (ii) in Theorem 4.1. Then it immediately follows from assertion (iii) in Theorem 4.1 that $\lim_{k\to\infty}\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}=0$.
@@
$\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}\leq$ $\tilde{\delta}c_{1}^{2}c_{3}\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}+(1-\tilde{\delta})\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F},$
where $c_{3}=\max(\sqrt{\mathcal{W}_{\mathbf{R}}})_{i_{1}i_{2}i_{3}}$. Then by assertion (iii) in Theorem 4.1, we have $\lim_{k\to\infty}\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}=0$.
@@
(19) $\begin{split}&H([X_{1}^{k}],[X_{2}^{k}],[X_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([X_{1}^{k+1}],[X_{2}^{k+1}],[X_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{4}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\end{split}$
@@
(20) $\left\|[H_{i}^{k}]^{\top}[{X}_{i}^{k}]-{[{X}_{i}^{k}]}^{\top}[H_{i}^{k}]\right\|_{F}\leq c_{5}\|[{X}_{i+1}^{k}]-[{X}_{i}^{k}]\|_{F},$
where $c_{4}=\frac{\alpha_{X}}{2(\alpha_{X}+2+\theta_{\max})^{2}}$, $\theta_{\max}=\max_{ijk}\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2}$, $c_{5}=2(\alpha_{X}\sqrt{n_{\max}}+\theta_{\max})$, $n_{\max}=\max\{n_{1},n_{2},n_{3}\}$, and
@@
To show the first inequality holds, the update of $[{X}_{i}^{k+1}]$ in (9) can be viewed as the iteration in (17) with $h([X])=\frac{1}{2}\|[X]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\|_{F}^{2}$, $[X]=[{X}_{i}^{k}]$, $[\bar{X}]=[{X}_{i}^{k+1}]$ and $\tau=\tilde{\alpha}_{X}=\frac{1}{1+\alpha_{X}}$. Then for $i=1$ we have
@@
where $c_{ij}^{k}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{X}}^{-1}+1+\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2})^{2}}\geq c_{4}$ and $\theta_{\max}$ is bounded, since the sequence $\{\mathcal{Z}^{k}\}$ is bounded. Using $[{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can obtain the fourth line above and rewrite part of the last line as follows
@@
$\begin{split}&H([{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{4}\left\|\left([\mathbf{I}_{m_{1}}]-[{X}_{1}^{k}]{[{X}_{1}^{k}]}^{\top}\right)[H_{1}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{4}\left\|\left([\mathbf{I}_{m_{2}}]-[{X}_{2}^{k}]{[{X}_{2}^{k}]}^{\top}\right)[H_{2}^{k}]\right\|_{F}^{2}\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{4}\left\|\left([\mathbf{I}_{m_{3}}]-[{X}_{3}^{k}]{[{X}_{3}^{k}]}^{\top}\right)[H_{3}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$[U^{k}][\Sigma^{k}][V^{k}]^{\top}=[{X}_{i}^{k}]-\tilde{\alpha}_{X}\left([{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)=\frac{1}{\alpha_{X}+1}\left(\alpha_{X}[X_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right),$
@@
(21) $[{X}_{i}^{k+1}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$ $=\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k+1}]$
$=(\alpha_{X}+1)[V]^{k}[\Sigma^{k}][V^{k}]^{\top}.$
@@
$=$ $\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k}]-[{X}_{i}^{k}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right)^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$-\left(\alpha_{X}[{X}_{i}^{k}]+[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right),$
where the last equation is obtained by (21). Taking the Frobenius norm of both sides, we obtain (20), where $2\|\alpha_{X}[{X}_{i}^{k}]^{(j)}+[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{F}\leq c_{5}$.
@@
Suppose that $\{\mathcal{Z}^{k}\}_{k\in\mathcal{K}}$ is a convergent subsequence of $\{\mathcal{Z}^{k}\}$ and converges to $\bar{\mathcal{Z}}$ as $k\in\mathcal{K}$ approaches $\infty$. By the updates of $\mathcal{S}^{k}$, $[\mathcal{G}^{k}]$ and $\mathcal{L}^{k}$, we have for any $k=0,1,\dots$
$\alpha_{\mathcal{S}}(\mathcal{S}^{k-1}-\mathcal{S}^{k})+$ $\delta(\mathcal{L}^{k}-\mathcal{L}^{k-1})$
$\in\delta(\mathcal{L}^{k}+\mathcal{S}^{k}-\mathcal{D})+\gamma\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\right)^{-1}\odot\partial\varphi(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}^{k}),$

### v2

@@ 3.2 The update of $[X_{i}]$
where $[P_{i}^{k}]=(\mathbf{R}(\mathcal{L}^{k}))_{[(i)]}$, $[Q_{1}^{k}]=\left([\mathcal{G}^{k}]\times_{2}[X^{k}_{2}]\times_{3}[X^{k}_{3}]\right)_{[(1)]}$, $[Q_{2}^{k}]=\big{(}[\mathcal{G}^{k}]\times_{1}[X_{1}^{k+1}]$$\times_{3}[X^{k}_{3}]\big{)}_{[(2)]}$, $[Q_{3}^{k}]=\left([\mathcal{G}^{k}]\times_{1}[X^{k+1}_{1}]\times_{2}[X^{k+1}_{2}]\right)_{[(3)]}$, and parameter $\tilde{\alpha}_{X}=\frac{\delta}{\delta+\alpha_{X}}$.
@@
Lemma 3.1 ([21, 2]).
Given $A\in\mathbb{R}^{m\times n}$, $m\geq n$, consider the following Stiefel manifold projection problem
@@ 3.3 The update of $[\mathcal{G}]$
(11) $[\mathcal{G}^{k+1}]=\operatorname{prox}_{\tilde{w}_{\mathcal{G}}\psi}\left([\mathcal{G}^{k}]-\tilde{\alpha}_{\mathcal{G}}([\mathcal{G}^{k}]-[\mathcal{O}^{k+1}])\right),$
where $[\mathcal{O}^{k+1}]=\mathbf{R}(\mathcal{L}^{k})\times_{1}[X_{1}^{k+1}]^{\top}\times_{2}[X_{2}^{k+1}]^{\top}\times_{3}[X_{3}^{k+1}]^{\top}$, $\tilde{w}_{\mathcal{G}}=\frac{1}{\delta+\alpha_{\mathcal{G}}}$, and $\tilde{\alpha}_{\mathcal{G}}=\frac{\delta}{\delta+\alpha_{\mathcal{G}}}$.
@@ 3.4 The update of $\mathcal{L}$
Before we compute $\mathcal{L}^{k+1}$, we define the transpose of $\mathbf{R}$ by $\mathbf{R}^{\top}:$ $\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ satisfying $\langle\mathbf{R}(\mathcal{L}),[\mathcal{Y}]\rangle=\langle\mathcal{L},\mathbf{R}^{\top}([\mathcal{Y}])\rangle$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, and $[\mathcal{Y}]\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$. Then we have
$\|\mathbf{R}(\mathcal{L})\|_{F}^{2}=\langle\mathcal{L},\mathbf{R}^{\top}\mathbf{R}(\mathcal{L})\rangle_{F}=\langle\mathcal{L},\mathcal{W}_{\mathbf{R}}\odot\mathcal{L}\rangle_{F},$
where $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}_{++}^{I_{1}\times I_{2}\times I_{3}}$ denotes a weight tensor such that the component-wise multiplication with $\mathcal{W}_{\mathbf{R}}$ is an equivalent operation of $\mathbf{R}^{\top}\mathbf{R}$, i.e., $\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\mathbf{R}^{\top}\mathbf{R}$, $\odot$ denotes the component-wise multiplication, and $\operatorname{Id}$ denotes the identity mapping on $\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$. Then $\mathcal{L}^{k+1}$ can be computed in a unique closed form as follows
(12) $\mathcal{L}^{k+1}=$ $(\delta\mathcal{W}_{\mathbf{R}}+\mathcal{I})^{-1}\odot\left(\delta\mathbf{R}^{\top}([\mathcal{Y}^{k+1}])+\mathcal{D}-\mathcal{S}^{k+1}\right),$
where $[\mathcal{Y}^{k+1}]$ denotes the approximated low-rank tensor
(13) $[\mathcal{Y}^{k+1}]=[\mathcal{G}^{k+1}]\times_{1}[X_{1}^{k+1}]\times_{2}[X_{2}^{k+1}]\times_{3}[X_{3}^{k+1}],$
$\mathcal{I}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ denotes the tensor with all entries equal to $1$, and $\mathcal{W}_{\mathbf{R}}^{-1}$ denotes the component-wise inverse of $\mathcal{W}_{\mathbf{R}}$, i.e., the $(i_{1},i_{2},i_{3})$-th entry of $\mathcal{W}_{\mathbf{R}}^{-1}$ is equal to $1/(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}$.
@@ 4 Convergence Analysis of The P-BCD Algorithm
${0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
@@
${0}\in\delta\nabla_{[\mathcal{G}]}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})+\partial\psi([\bar{\mathcal{G}}]),$
${0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\nabla_{\mathcal{L}}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),$
@@
(15a) ${0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
@@
(15e) ${0}\in\delta([\bar{\mathcal{G}}]-[\bar{\mathcal{O}}])+\partial\psi([\bar{\mathcal{G}}]),$
(15f) ${0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\mathcal{W}_{\mathbf{R}}\odot\bar{\mathcal{L}}-\delta\mathbf{R}^{\top}([\bar{\mathcal{Y}}]),$
@@
$\geq$ $\frac{\alpha_{\mathcal{S}}}{2}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F},$
where $\alpha>0$. Next, it follows from the update of $[X_{i}]$ and Lemma 3.1 that
@@
Combining the inequalities above, we obtain (16) with $\alpha=\min\{\alpha_{\mathcal{S}},\alpha_{\mathcal{G}},\alpha_{X}\}>0$.
(ii) Since $[X_{i}^{k}]^{\top}[X_{i}^{k}]=[\mathbf{I}_{n_{i}}]$ for each $i=1,2,3$, we have the sequence $\{[X_{i}^{k}]\}$ is bounded. By (i), we have $\Phi(\mathcal{Z}^{k})\leq\Phi(\mathcal{Z}^{0})$. Also, we observe that $\Phi(\mathcal{Z}^{k})\geq\gamma\varphi(\mathcal{S}^{k})+\psi([\mathcal{G}^{k}])\geq 0$. Since
@@
$\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}\leq$ $\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}+c_{1}\|[{X}_{1}^{k+1}]-[{X}_{1}^{k}]\|_{F}$
$+c_{1}\|[{X}_{2}^{k+1}]-[{X}_{2}^{k}]\|_{F}+c_{1}\|[{X}_{3}^{k+1}]-[{X}_{3}^{k}]\|_{F},$
where $c_{1}=\max_{k}\|[\mathcal{G}^{k}]\|_{F}<\infty$ according to assertion (ii) in Theorem 4.1. Then it immediately follows from assertion (iii) in Theorem 4.1 that $\lim_{k\to\infty}\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}=0$.
@@
$\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}\leq$ $\delta\sqrt{c_{2}}\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}+\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F},$
where $c_{2}=\max(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}$. Then by assertion (iii) in Theorem 4.1, we can have $\lim_{k\to\infty}\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}=0$.
@@
(19) $\begin{split}&H([X_{1}^{k}],[X_{2}^{k}],[X_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([X_{1}^{k+1}],[X_{2}^{k+1}],[X_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\end{split}$
@@
(20) $\left\|[H_{i}^{k}]^{\top}[{X}_{i}^{k}]-{[{X}_{i}^{k}]}^{\top}[H_{i}^{k}]\right\|_{F}\leq c_{4}\|[{X}_{i+1}^{k}]-[{X}_{i}^{k}]\|_{F},$
where $c_{3}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{X}}^{-1}+1+\theta_{\max})^{2}}>0$, $c_{4}=2(\alpha_{X}\sqrt{n_{\max}}+\theta_{\max})>0$, $\theta_{\max}=\max_{ijk}$ $\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2}$, $n_{\max}=\max\{n_{1},n_{2},n_{3}\}$, and
@@
To show the first inequality holds, the update of $[{X}_{i}^{k+1}]$ in (9) can be viewed as the iteration in (17) with $h([X])=\frac{1}{2}\|[X]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\|_{F}^{2}$, $[X]=[{X}_{i}^{k}]$, $[\bar{X}]=[{X}_{i}^{k+1}]$ and $\tau=\tilde{\alpha}_{X}=\frac{\delta}{\delta+\alpha_{X}}$. Then for $i=1$ we have
@@
where $c_{ij}^{k}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{X}}^{-1}+1+\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2})^{2}}\geq c_{3}>0$ and $\theta_{\max}$ is bounded, since the sequence $\{\mathcal{Z}^{k}\}$ is bounded. Using $[{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can obtain the fourth line above and rewrite part of the last line as follows
@@
$\begin{split}&H([{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{1}}]-[{X}_{1}^{k}]{[{X}_{1}^{k}]}^{\top}\right)[H_{1}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{2}}]-[{X}_{2}^{k}]{[{X}_{2}^{k}]}^{\top}\right)[H_{2}^{k}]\right\|_{F}^{2}\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{3}}]-[{X}_{3}^{k}]{[{X}_{3}^{k}]}^{\top}\right)[H_{3}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$[U^{k}][\Sigma^{k}][V^{k}]^{\top}=[{X}_{i}^{k}]-\tilde{\alpha}_{X}\left([{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)=\frac{1}{\alpha_{X}+\delta}\left(\alpha_{X}[X_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right),$
@@
(21) $[{X}_{i}^{k+1}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$ $=\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k+1}]$
$=(\alpha_{X}+\delta)[V]^{k}[\Sigma^{k}][V^{k}]^{\top}.$
@@
$=$ $\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k}]-[{X}_{i}^{k}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right)^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$-\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right),$
where the last equation is obtained by (21). Taking the Frobenius norm of both sides, we obtain (20), where $2\|\alpha_{X}[{X}_{i}^{k}]^{(j)}+\delta[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{F}\leq c_{4}$.
@@
Suppose that $\{\mathcal{Z}^{\tilde{k}}\}_{\tilde{k}\in\mathcal{K}}$ is a convergent subsequence of $\{\mathcal{Z}^{k}\}$ and converges to $\bar{\mathcal{Z}}$ as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. By the updates of $\mathcal{S}^{\tilde{k}}$, $[\mathcal{G}^{\tilde{k}}]$ and $\mathcal{L}^{\tilde{k}}$, we have for any $\tilde{k}\in\mathcal{K}$,
$\alpha_{\mathcal{S}}(\mathcal{S}^{\tilde{k}-1}-\mathcal{S}^{\tilde{k}})+$ $(\mathcal{L}^{\tilde{k}}-\mathcal{L}^{\tilde{k}-1})$
$\in\mathcal{L}^{\tilde{k}}+\mathcal{S}^{\tilde{k}}-\mathcal{D}+\gamma\partial\varphi(\mathcal{S}^{\tilde{k}}),$